OpenStudy (anonymous):

Let f(x) = x4 – 8x3 + 16x2 – 19 and g(x) = x – 5.

OpenStudy (saifoo.khan):

And..?

OpenStudy (anonymous):

find f(x) over g (x)

Parth (parthkohli):

Can you find this: \[{x^4 - 8x^3 + 16x^2 - 19 \over x -5} \]

OpenStudy (anonymous):

could i use synthetic division...

OpenStudy (saifoo.khan):

Yes you can/

Parth (parthkohli):

Of course! I recommend it.

OpenStudy (anonymous):

ok. let me work this out real quick

OpenStudy (anonymous):

\[1^3 - 3^2+1 R=-14\]

OpenStudy (anonymous):

the remainder is negative 14

Parth (parthkohli):

Wait a second; I'd do that on my paper.

OpenStudy (anonymous):

k

Parth (parthkohli):

I find the remainder to be 6.

OpenStudy (anonymous):

wait of my question or the one u gave me

Parth (parthkohli):

Your question is just the same as my question.

OpenStudy (anonymous):

lol was not payin attention .... could explain your answer

Parth (parthkohli):

All right. Given that \(f(x) = x^4 - 8x^3 + 16x^2 - 19\) and \(g(x) = x - 5\), we may say: \[{f(x) \over g(x)} = {x^4 - 8x^3 + 16x^2 - 19 \over x -5} \] Does that make sense?

OpenStudy (anonymous):

yes

Parth (parthkohli):

All right, so can you find that?

Parth (parthkohli):

By finding, I meant simplifying.

OpenStudy (anonymous):

yea i got it now i forgot to follow the f(x) over g(x)......